OpenStudy (anonymous):

if f(x)=2x-3 & g(x)=x What is f(g(2))?

OpenStudy (anonymous):

A) -2 B) 0 C) 1 D) 4

OpenStudy (anonymous):

@Fall12-13

OpenStudy (anonymous):

F(x) = 2x - 3 G(x) = x F(G(x)) = F(x) = 2x - 3 F(G(2)) = F(2) = 2(2) - 3 = 4 - 3 =1 Which is C.

OpenStudy (anonymous):

thank you(: